A force was applied to car whose mass is 2,100 kg. The car then accelerates at 0.7 m/s2. If the same force is applied to another car whose mass is 100 kg lesser, what is the resulting acceleration? [Hint: Solve for the value of the force applied to the 1st car. Then, compute the mass of the 2nd car, and solve for its acceleration.]


1
Expert's answer
2021-11-02T10:00:09-0400

From the question, the amount of force acting on the two cars is same. So,

"F=m_1a_1=m_2a_2"

The acceleration of second car is

"a_2=(\frac{m_1}{m_2})a_1=(\frac{100 kg}{2100 kg})(0.7 m/s^2)=0.33 m/s^2"